2. Suppose you take a multiple choice test with 10 questions, and each question has 5 answer choices (a. (10 points) b, c.d, e). What is the probability you get exactly 4 questions correct just by guessing? on)
Answer #1

P(x) = nCx px qn-x

Probability of correct answer, p = 1/5 = 0.2

q = 1 - p = 0.8

n = Number of trials(questions) = 10

p(getting exactly 4 questions correct) = P(4)

= 10C4 x 0.24 x 0.86

= 0.0881